OpenStudy (anonymous):

n over 56=5 over 40 A. n = 8 B. n = 7 C. n = 35 D. n = 240

OpenStudy (anonymous):

Its just an eqn, n/56=5/40 Now find for n, by taking 56 to other side

OpenStudy (anonymous):

wats the answer

OpenStudy (anonymous):

I cant say it. Just do the calculation n=5*56/40

OpenStudy (anonymous):

240

OpenStudy (anonymous):

u got it wrong

OpenStudy (anonymous):

oh welll...